OpenStudy (anonymous):

Simplify the expression. (-1^5/-2^-3)^2 A. 1/64 B. 64 C. 2^30 D. 2^-30

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

.1^5=-1 -2^-3= -1/8 -1/-1/8=1/8^2 1/64

OpenStudy (anonymous):

thanks can you help with more if I need help?

OpenStudy (anonymous):

sure